A​ BBB-rated corporate bond has a yield to maturity of

6.3 %

A U.S. treasury security has a yield to maturity of

4.3 %

These yields are quoted as

APRs

with semiannual compounding. Both bonds pay​ semi-annual coupons at a rate of

5.0 %

and have five years to maturity.    

a. What is the price​ (expressed as a percentage of the face​ value) of the treasury​ bond?

b. What is the price​ (expressed as a percentage of the face​ value) of the​ BBB-rated corporate​ bond?

c. What is the credit spread on the BBB​ bonds?
Answer #1

Solution:

a)

In order to calculate the price of the Tbond we will be using BAII plus calcutor TVM row

Enter the given values as

FV=100 , I?Y =4.3 (annually) for semi annual we have 6.3/2 = 2.15, Coupon = ( 5% of 100 ) /2 = 2.5 , N = 5*2 =10,Pv=?

After entering the values we Press CPT+PV, PV = 103.119 or 103.119 % as a percentage of par

b) In order to calculate the price of the BBBbond we will be using BAII plus calcutor TVM row

Enter the given values as

FV=100 , I?Y =6.3 (annually) for semi annual we have 6.3/2 = 3.15, Coupon = ( 5% of 100 ) /2 = 2.5 , N = 5*2 =10,Pv=?

After entering the values we Press CPT+PV, PV = 94.49 or 94.49% as a percentage of par

c) Credit spread on a risky bond

Yield of a risky bond - Yiled of a Tbond

= 6.3 - 4.3 = 2% or 200 Basis point is the credit spread